the sum and product of slopes of tangents to parabola y^2=4ax drawn from the point (2,_3) are
 

Grade:12th pass

1 Answers

Vikas TU
6 years ago
The digression at parabola at (at12,2at1) is yt1 =x+at12 - - 1 
so digression at (at22,2at2) is yt2 =x+at22 - - 2 
Give these digression a chance to meet at point P [ Here point P is (2,- 3) 
So the point p can be spoken to as {at1t2,a(t1 +t2)} 
Thus, at1t2 =2 and a(t1 +t2) = - 3 
a(t1 +t2) = - 3 
at1 + at2 = - 3 
2/t2 + at2 = - 3 
at2 ^2 +2 = - 3t2 
The estimation of incline can't be found until the estimation of an is given .Hence the question is inadequate.